# Machine learning applications in archaeological practices: a review

### v2

@@ 3.3.1 Statistical overview
Figure 4: Number of articles published per country based on the country of the first author’s affiliation.
Examining the different architecture of machine learning methods used, we observed that artificial neural networks (ANNs) are the most frequently used with 110 occurrences across all study cases, followed by ensemble learning with 70 occurrences, together representing 62% of all machine learning applications study cases in our review [Figs. 2 and 5B]. Linear classifiers, decision trees, rule induction, and nearest neighbour classifiers were also well-represented, accounting for 25% of total applications. An increase of the use of ANNs in our list of reviewed articles from 2021 onwards could be observed, with the number of applications multiplying four-fold compared to the number of applications in the years 2019 and 2020 [Fig. 5B]. At a more granular scale, we found a total of 70 machine learning models [Supplementary file]. In addition, we found a particularly high diversity in the unsupervised learning and clustering category of model families, with a unique model for each of the twelve recorded applications for the category [Supplementary file]. If ANNs are the most represented model family, the most common model was random forest, with 54 study cases counted. We found that across all articles reviewed, the mean number of models used per study case was 2.12, with a high disparity across different publications, with some applying only one model, while articles such as Bataille et al. [2018] and Courtenay et al. [2019] tested six different models at once. Classification was the most common form of data evaluation, applied in more than 75% of study cases (n=112). On the other hand, only 19% (n=29) were regression models, and clustering was only represented with six uses [6%, Figs. 2, 7, 8 and 9].
It was more difficult to observe patterns for the other features we collected, as the attribution to one or several categories for an article was more subjective to the person reviewing. Two of the a posteriori study tasks are highly prominent: automatic structure detection and artefact classification. These two tasks account for 45% (n=67) of all study case tasks [Fig. 2]. Tasks such as taphonomic classification, archaeological predictive models, and architectural element classification are well represented with 5 to 13 study cases each, whilst nine task categories (e.g. sourcing, species classification, movement recognition) were only relevant for around 5 study cases. No clear pattern could be distinguished diachronically for how the desired scientific goal of machine learning applications have evolved over time.
@@
Finally, according to the assessment of article authors or (if not discussed) the results metrics, we could observe that the majority of applications (79) had reported “successful” outcomes [Fig. 6B], with an additional 40 mixed results (27.2%). Only 15 applications (10.2%) were reported to have been “unsuccessful”, whilst we noted 9 (6.1%) study cases whose application had important methodological issues, and classified them as such even if the authors reported a successful outcome. Finally, four (2.7%) study cases did not have a defined outcome. These results, however, could be affected by the well-reported issue of publication bias against the reporting of negative results (Song et al., 2013), as well as the related file-drawer-effect (Rosenthal, 1979; Sparks, 2009), which can strongly affect the perception and production of scientific results (Song et al., 2010).
Figure 5: (A) Number of articles from each archaeological subfield between 1997 and 2023. (B) Number of articles from each architecture class between 1997 and 2023. Empty bar charts represent the number 1.
@@ 3.3.2 Aggregated information
Correlations between tasks and model families are highly heterogeneous [Fig. 7]. Whilst the highest correspondence rate between taphonomic classification and unsupervised learning and clustering methods is 26%, the correspondence between artefact classification and ANNs is much higher, at 49%. Even stronger correlation rates could be observed between archaeological predictive models and ensemble learning methods at 66%, and between automatic structure detection and ANNs, at 60%.
The correlations between archaeological subfields and study tasks are lower than the correlations visible for architecture [Fig. 8]. Articles dealing with environmental reconstruction had their highest correlation at only 33% with the subfield of geoarchaeology. Studies treating artefact classification problematics have a higher 45% correspondence rate with subfields of classification and typology. The highest correlation rate, however, is present between the task of automatic structure detection and the subfield of surveying, at 80% correlation.
@@
Figure 6: (A) The five more represent classes of input data among the reviewed papers, n = 148. (B) Results of the reviewed papers according to the authors or presented results, n = 147.
@@ 3.3.3 Classification of artefacts and animal remains
Whilst classification tasks were mainly performed on ceramic material (31%), as in Hörr et al. [2014] or Anichini et al. [2021], a wide range of different materials were used in our dataset of reviewed articles. Studies on coins (Boon et al., 2009), stone tools (MacLeod, 2018; Pargeter et al., 2019; Emmitt et al., 2022), archaeobotanical seed remains (Landa et al., 2021), phytoliths (Berganzo-Besga et al., 2022), ivory figurines (Gansell et al., 2014), were observed. Moreover, 17% of study cases were conducted on cave art images, such as in Kogou et al. [2020], or Horn et al. [2022a; 2022b]. On all classification study cases 17% (n=5) dealt with the bioarchaeological classification of human or animal remains. As a consequence of the variety of materials studied, the input data for models were very diverse. Small-scale images accounted for 48% (n=14) of all input data for classification task, metric data for 27% (n=8), as well as less frequent input data types such as multilayer images of artefact (e.g. depth maps, 3D model layers), spectra and remote sensing images all together accounted for 23% (n=7).
@@ 3.3.4 Archaeological predictive models (APMs)
Ten studies in our review focused on archaeological predictive models (APMs), dealing with either human occupation, or human-environment interactions. Input data were in 77% (n=7) of all study case larger-scale raster images from a diverse range of natural covariates or remote sensing images.
@@
Figure 7: Alluvial diagram of the different tasks in the analysed studies on the left, the related architecture of machine learning models on the right and the evaluation process in the background. Tasks and architectures poorly represented (n < 5) have been classified as “others”. A study might have applied numerous models or its research objectives could be classified into more than one task. In such cases, we created multiple entries for each paper where applicable [Supplementary file].
@@ 3.3.5 Automatic structures detection
Representing 25% (n=38) of all study cases, automatic structures detection, also known as geographical object-based image analysis (GEOBIA), is the most prominent task category in our review. Input data came from a wide range of remote sensing images, mainly LiDAR or airborne laser scanning (ALS), used in 50% of all study cases in this task, or various satellite image collections (23%), such as in Menze et al. [2006] or Orengo et al. [2020]. The remaining 27% input data used were either UAV ortho-images, such as in Orengo and Garcia-Molsosa [2019], Monna et al. [2020], Agapiou et al. [2021], Altaweel et al. [2022], and Fisher et al. [2022], historical maps, such as Garcia-Molsosa et al. [2021], and since 2020, ground-penetrating radar (GPR), or even bathymetric data (Febriawan et al., 2020; Bordon et al., 2021). ANNs were the most prominent family of methods for automatic structure detection, accounting for 62% of all models applied for this task. Different ANN models were used, the most popular being mask region-based convolutional neural networks (MR-CNN), which accounted for 13% of all models applied for automatic structure detection, followed by faster region-based convolutional neural networks (FR-CNN), and U-Net, which each accounted for 8% and 7% of the models used. However, the most popular model was random forest, which accounted for 16% of all models used for this task. The results of applications in this task were mainly reported as unsuccessful or partially successful (59%, n=31).
Figure 8: Alluvial diagram of the different tasks in the analysed studies on the left, the related archaeological subfields on the right with the evaluation process in the background. Tasks and subfields poorly represented (n < 5) have been classified as “others”. A study might have been attributed to several subfields or its research objectives could be classified into more than one task. In such cases, we created multiple entries for each paper where applicable [Supplementary file].
@@ 3.3.8 Taphonomic classification
Figure 9: Alluvial diagram of the different tasks in the analysed studies on the left, the related results on the right with the evaluation process in the background. Tasks poorly represented (n < 5) have been classified as “others”. A study might have its research objectives classified into more than one task. In such cases, we created multiple entries for each paper where applicable [Supplementary file].
@@ 4.1 General considerations on the corpus
A possible initial observation is the comparatively late interest in machine learning applications in the field of archaeology compared to other sciences (Dramsch, 2020; Padarian et al., 2020; Shehab et al., 2022; Wang et al., 2022). Indeed, the exponential increase in publications related to machine learning applications in archaeology can be dated from the period between 2018 and 2019 onwards [Fig. 3]. This late development could arguably be attributed to a tradition of suspicion in archaeology (Yoffee and Fowles, 2010). Researchers sometimes consider archaeological data as scattered, complex (Smith and Peregrine, 2011, p.7) and unable to be explained by models that are not, fully, human-made. The uncertainty for selecting the best model to analyse the complexity of past human action is not recent (Dorans, 1990) but has been exacerbated with machine learning and even more with deep learning models, which are often black boxes, difficult to interpret the reasoning behind their choice and weighing of features (Ramazzotti, 2020; Fiorucci et al., 2020; Bickler, 2021; Calder et al., 2022).
@@
Regarding publication policies across our list of reviewed articles, the high representation of open-access papers in a field where the traditional closed-access model of publications is widespread [see below; Marwick and Birch 2018] could be a signal for a paradigm change in which scholars seek to share their knowledge more broadly and fairly (Marwick et al., 2017; Nicholson et al., 2023). While many journals that deal specifically with computer science applications in archaeology (e.g. Archaeologia e Calculatori, Internet Archaeology, or Journal of Computer Applications in Archaeology) only have open access publication routes, more classical and broader journals like the Journal of Archaeological Research, Journal of Archaeological Science or Antiquity provide mainly traditional non-open access papers, with open access papers comprising only (respectively) 10%, less than 5% and less than 2% of all articles on the 1st of September 2024. These results are unlike observations made for articles dealing with machine learning applications in other disciplines such as soil science (Padarian et al., 2020), where paid-access routes vastly dominate that field of research. However, researchers have to keep in mind the existence of non-open access journal articles and book chapters to avoid a FUTON bias (Wentz, 2002), which will lead to not exploiting part of the wider research results in the field. This points to a wider adoption of both FAIR (Findable, Accessible, Interoperable, Reusable) data principles (Wilkinson et al., 2016; Nicholson et al., 2023)(Wilkinson et al. 2016; Nicholson et al. 2023) and CARE (Collective benefit, Authority to control, Responsibility, Ethics) principles (Carroll et al., 2020; Gupta et al., 2023)(Carroll et al. 2020; Gupta et al. 2023). Nonetheless, if the degree of open access of machine learning applications in archaeology seems promising, the large disparity borne from colonialism remains unfortunately wide (Davis, 2020). The difference in the number of publications from “Global North” and “Global South” countries can be directly linked with the gap in the economic development between higher and lower-income countries, leading to the underdevelopment of science and technology (Sagasti, 1973; Allik et al., 2020). Increasing collaborations between different countries and institutions could fill this gap between higher and lower-income countries (Sonnenwald, 2007).
@@ 4.2 How is machine learning used across archaeology?
In the earlier stage of machine learning applications in archaeology, many archaeological questions could only be answered by seeking experts in informatics, mathematics, and physics (Menze et al., 2006; Boon et al., 2009; Menze and Ur, 2012; Hörr et al., 2014). However a noticeable shift has occurred over time. Archaeologists have progressively embraced programming their own tools (e.g. Python, R, Matlab, Java), and have also turned to specialised software solutions such as Grilli and Remondino [2019], Garcia-Malsosa et al. [2021], Demjam et al. [2022] or Casini et al. [2023]. Nowadays, in many studies, archaeologists are both asking and answering the questions themselves (Orengo and Garcia-Molsosa, 2019; Orengo et al., 2020; Yaworsky et al., 2020; Carter et al., 2021). The CAA Special Interest Group on Scientific Scripting Languages in Archaeology (SIG-SSLA ) is one example of this polyvalency of archaeologists developing scripting or machine learning tools for their problems. In certain instances, computer science specialists independently pose questions and generate tailored answers for specific applications (Toler-Franklin et al., 2010; Ushizima et al., 2020; Wunderlich et al., 2022). However, there is always a question of the ratio between the time spent developing new skills (Aldenderfer, 1998, pp. 109-112) and the increasing complexity, if not in programming, then in understanding, of machine learning methods, making it more and more difficult for an archaeologist to master all abilities needed to develop an entire workflow for a computer-based approach.
@@ 4.2.1 Classification of artefacts and animal remains
Classifying artefacts and animal (including human) remains into categories is a central aspect of archaeological work (Read, 2018). Archaeologists have looked for solutions in computer science since the second half of the twentieth century [cf. 1 Binford and Binford 1966, Orton 1980, pp. 156-178]. It is therefore not surprising to see machine learning applications developed for this purpose already very early [Fig. 5A]. Earlier papers generally applied methods such as decision trees, ensemble learning, Bayesian classifiers, as well as linear classifiers combined with nearest neighbour classifiers (Nguifo et al., 1997; Hörr et al., 2014; Mircea et al., 2015; MacLeod, 2019; Canul-Ku et al., 2019; Pargeter et al., 2019; González-Molina et al., 2020). Since 2020, however, ANNs have gained more traction [e.g. Ushizima et al. 2020; Graham et al. 2020], and now represent the most commonly used family of methods applied for this task [Fig. 7]. This choice is reasonable, as ANNs, and convolutional neural networks in particular, are at the forefront for computer vision-based tasks such as the classification of objects in images (Guo et al., 2016; Chai et al., 2021). Dimensionality reduction methods (e.g. PCA and MDA) are often performed alongside machine learning algorithms, and can be an effective way to improve model accuracy when applied correctly (e.g. Muzzall 2021).
@@
Concerns about the validity of certain forms of typological classification, both human-made and computer-based, as a practice have been previously raised, especially from North American researchers (Adams and Adams, 1991; Hörr et al., 2014). Other researchers, however, do not hesitate in using computational approaches for classification (Pargeter et al., 2019), and even creating new classification methodologies with the help of machine learning. Notably, Hörr and others, seeking a less subjective approach to typological classification, developed a three-layer method for this purpose [Hörr 2011, pp. 136-137, Hörr et al. 2014], successively incorporating unsupervised, semi-supervised, and supervised classification algorithms within a framework Fayyad et al. [1996] termed the Knowledge Discovery Process. This approach is promising for the analysis of archaeological material, and could prove useful for many archaeological studies. Such an approach invites critical questions on the artificiality of typological classifications and opens new avenues for future studies such as those dealing with ceramic material or stone tools. We can also mention Martín-Perea et al. [2020], who used a method similar to that of (Hörr et al., 2014) with a three-level pipeline to detect fossiliferous levels of an archaeological or palaeontological site.
@@
Assessing and comparing the outcomes of these diverse classifications is challenging, given the inherent variations in datasets, modelling techniques, and evaluation metrics employed across different studies. Nonetheless, various opinions have already emerged on the issue. Emmit et al. [2022] and Jalandoni [2022] are optimistic about using machine learning for the classification of archaeological material, though we could also add Hörr et al. [2014], who are more cautious but remain hopeful for the future of the practice. Others are less enthusiastic, however, and point to mixed results (Demjan et al., 2022; Lyons et al., 2022). Overall it seems likely that machine learning applications for the classification of archaeological material will spread not only within the scientific community but also among professionals of archaeology, conservation institutions, and a wider public. The ArchAIDE project is already a noteworthy example of an application of open-access machine learning tools for archaeology to serve the public at large (Gualandi et al., 2021; Anichini et al., 2021).

### v3

@@ 3.3.1 Statistical overview
Figure 4: Number of articles published per country based on the country of the first author’s affiliation. Figure generated with R 4.2.2 (code available in supplementary material 3).
Examining the different architecture of machine learning methods used, we observed that artificial neural networks (ANNs) and ensemble learning represent 62% of all study cases, with 110 and 70 occurrences respectivly [Figs. 2 and 5B]. Linear classifiers, decision trees, rule induction, and nearest neighbour classifiers were also well-represented, accounting for 25% of total applications. An increase of the use of ANNs in our list of reviewed articles from 2021 onwards could be observed, with the number of applications multiplying four-fold compared to the number of applications in the years 2019 and 2020 [Fig. 5B]. At a more granular scale, we found a total of 70 machine learning models [supplementary file]. In addition, we found a particularly high diversity in the unsupervised learning and clustering category of model families, with a unique model for each of the twelve recorded applications for the category [supplementary file]. Despite the dominance of ANNs, random forest was the most common individual model (n = 54). We found that across all articles reviewed, the mean number of models used per study case was 2.12, with a high disparity across different publications, with some applying only one model, while articles such as Bataille et al. [2018] and Courtenay et al. [2019] tested six different models at once. Classification was the most common form of data evaluation, applied in more than 75% of study cases (n = 112). On the other hand, only 19% (n = 29) were regression models, and clustering was only represented with six uses [6%, Figs. 2, 7, 8 and 9].
It was more difficult to observe patterns for the other features we collected, as the attribution to one or several categories for an article was more subjective to the person reviewing. Two of the a posteriori study tasks are highly prominent: automatic structure detection and artefact classification. These two tasks account for 45% (n = 67) of all study case tasks [Fig. 2]. Tasks such as taphonomic classification, archaeological predictive models, and architectural element classification are well represented with 5 to 13 study cases each, whilst nine task categories (e.g. sourcing, species classification, movement recognition) were only relevant for around 5 study cases.
@@
Finally, according to the assessment of article authors or (if not discussed) the results metrics, we could observe that most applications (n = 79) had reported “successful” outcomes [Fig. 6B], with an additional 40 mixed results (27.2%). Only 15 applications (10.2%) were reported to have been “unsuccessful”, whilst we noted 9 (6.1%) study cases whose application had important methodological issues, and classified them as such even if the authors reported a successful outcome. Finally, four (2.7%) study cases did not have a defined outcome. These results, however, could be affected by the well-reported issue of publication bias against the reporting of negative results (Song et al., 2013), as well as the related file-drawer-effect (Rosenthal, 1979; Sparks, 2009), which can strongly affect the perception and production of scientific results (Song et al., 2010).
Figure 5: Fig. 5 (A) Number of articles from each archaeological subfield between 1997 and 2023. (B) Number of articles from each architecture class between 1997 and 2023. Empty bar charts represent the number 1. Figure generated with R 4.2.2 (code available in supplementary material 3).
@@ 3.3.2 Aggregated information
Correlations between tasks and model families are highly heterogeneous [Fig. 7]. Whilst the highest correspondence rate between taphonomic classification and unsupervised learning and clustering methods is 26%, the correspondence between artefact classification and ANNs is much higher, at 49%. Even stronger correlation rates could be observed between archaeological predictive models and ensemble learning methods at 66 and between automatic structure detection and ANNs, at 60%.
The correlations between archaeological subfields and study tasks are lower than the correlations visible for architecture [Fig. 8]. Articles dealing with environmental reconstruction had their highest correlation at only 33 with the subfield of geoarchaeology. Studies treating artefact classification problematics have a higher 45% correspondence rate with subfields of classification and typology. The highest correlation rate, however, is present between the task of automatic structure detection and the subfield of surveying, at 80% correlation.
@@
Figure 6: (A) The five more represent classes of input data among the reviewed papers, n = 148. (B) Results of the reviewed papers according to the authors or presented results, n = 147. Figure generated with R 4.2.2 (code available in supplementary material 3).
@@ 3.3.3 Classification of artefacts and animal remains
Whilst classification tasks were mainly performed on ceramic material (31%), as in Hörr et al. [2014] or Anichini et al. [2021], a wide range of different materials were used in our dataset of reviewed articles. Studies on coins (Boon et al., 2009), stone tools (MacLeod, 2018; Pargeter et al., 2019; Emmitt et al., 2022), archaeobotanical seed remains (Landa et al., 2021), phytoliths (Berganzo-Besga et al., 2022), ivory figurines (Gansell et al., 2014), were observed. Moreover, 17% of study cases were conducted on cave art images, such as in Kogou et al. [2020], or Horn et al. [2022a; 2022b]. On all classification study cases 17% (n = 5) dealt with the bioarchaeological classification of human or animal remains. Due to the variety of materials studied, the input data for models were very diverse. Small-scale images accounted for 48% (n = 14) of all input data for classification task, metric data for 27% (n = 8), as well as less frequent input data types such as multilayer images of artefact (e.g. depth maps, 3D model layers), spectra and remote sensing images all together accounted for 23% (n = 7).
@@ 3.3.4 Archaeological predictive models (APMs)
Ten studies in our review focused on archaeological predictive models (APMs), dealing with either human landscape occupation patterns, or human-environment interactions. Input data were in 77% (n = 7) of all study case larger-scale raster images from a diverse range of natural covariates or remote sensing images.
@@
Figure 7: Alluvial diagram of the different tasks in the analysed studies on the left, the related architecture of machine learning models on the right and the evaluation process in the background. Tasks and architectures poorly represented (n < 5) have been classified as “others”. A study might have applied numerous models or its research objectives could be classified into more than one task. In such cases, we created multiple entries for each paper where applicable [supplementary file]. Figure generated with R 4.2.2 (code available in supplementary material 3).
@@ 3.3.5 Automatic structures detection
Representing 25% (n = 38) of all study cases, automatic structures detection, also known as geographical object-based image analysis (GEOBIA), is the most prominent task category in our review. Input data came from a wide range of remote sensing images, mainly LiDAR or airborne laser scanning (ALS), used in 50% of all study cases in this task, or various satellite image collections (23%), such as in Menze et al. [2006] or Orengo et al. [2020]. The remaining 27% input data used were either UAV ortho-images, such as in Orengo and Garcia-Molsosa [2019], Monna et al. [2020], Agapiou et al. [2021], Altaweel et al. [2022], and Fisher et al. [2022], historical maps, such as Garcia-Molsosa et al. [2021], and since 2020, ground-penetrating radar (GPR), or even bathymetric data (Febriawan et al., 2020; Bordon et al., 2021).
ANNs were the most prominent family of methods for automatic structure detection, accounting for 62% of all models applied for this task. Different ANN models were used, the most popular being mask region-based convolutional neural networks (MR-CNN), which accounted for 13% of all models applied for automatic structure detection, followed by faster region-based convolutional neural networks (FR-CNN), and U-Net, which each accounted for 8% and 7% of the models used. However, the most popular model was random forest, which accounted for 16% of all models used for this task. The results of applications in this task were mainly reported as unsuccessful or partially successful (59%, n = 31).
Figure 8: Alluvial diagram of the different tasks in the analysed studies on the left, the related archaeological subfields on the right with the evaluation process in the background. Tasks and subfields poorly represented (n < 5) have been classified as “others”. A study might have been attributed to several subfields or its research objectives could be classified into more than one task. In such cases, we created multiple entries for each paper where applicable [supplementary file]. Figure generated with R 4.2.2 (code available in supplementary material 3).
@@ 3.3.8 Taphonomic classification
Many of the articles reviewed applied a large number of machine learning methods as well as methods that we do not consider machine learning [such as PLSDA and MDA; e.g. Domínguez-Rodrigo 2018; Abellán et al. 2022] to compare their performance on the specific task, whilst others [e.g. Byeon et al. 2019; Cifuentes-Alcobendas and Domínguez-Rodrigo 2019] were more restrained. This meant that there was no clear machine learning method or family of methods that was used much more frequently than others, although neural networks were applied in all but one of the articles reviewed (see “Results”), which is not the case for other methods also commonly used in these articles, such as SVMs or random forests.
@@
Figure 9: Alluvial diagram of the different tasks in the analysed studies on the left, the related results on the right with the evaluation process in the background. Tasks poorly represented (n < 5) have been classified as “others”. A study might have its research objectives classified into more than one task. In such cases, we created multiple entries for each paper where applicable [supplementary file]. Figure generated with R 4.2.2 (code available in supplementary material 3).
@@ 4.1 General considerations on the corpus
A possible initial observation is the comparatively late interest in machine learning applications in the field of archaeology compared to other sciences (Dramsch, 2020; Padarian et al., 2020; Shehab et al., 2022; Wang et al., 2022). Indeed, the exponential increase in publications related to machine learning applications in archaeology can be dated from the period between 2018 and 2019 onwards [Fig. 3]. This late development could arguably be attributed to a tradition of suspicion in archaeology (Yoffee and Fowles, 2010). Researchers sometimes consider archaeological data as scattered, complex (Smith and Peregrine, 2011, p.7) and unable to be explained by models that are not, fully, human-made. The uncertainty for selecting the best model to analyse the complexity of past human action is not recent (Doran, 1990) but has been exacerbated with machine learning and even more with deep learning models, often opaque, “black boxes” in their internal logic (Ramazzotti, 2020; Fiorucci et al., 2020; Bickler, 2021; Calder et al., 2022). Furthermore, it appears that some sub-disciplines of archaeology have until 2022 yet to exploit the new development of machine learning to the same degree as other sub-disciplines. Following our classification of archaeological subfields (see supplementary material), ethnoarchaeology, archaeobotany, and cognitive archaeology are under-represented in our corpus [Figs. 2 5B]. Although the reasons for such an underrepresentation remain unclear, it is likely that many research questions in these subfields are less suitable for machine learning applications at this current time, or at least to lesser degree than other sub-disciplines (Marom, 2025).
@@
Regarding publication policies across our list of reviewed articles, the high representation of open-access papers in a field where the traditional closed-access model of publications is widespread [see below; Marwick and Birch 2018] could be a signal for a paradigm change in which scholars seek to share their knowledge more broadly and fairly (Marwick et al., 2017; Nicholson et al., 2023). While many journals that deal specifically with computer science applications in archaeology (e.g. Archaeologia e Calculatori, Internet Archaeology, or Journal of Computer Applications in Archaeology) only have open access publication routes, more classical and broader journals like the Journal of Archaeological Research, Journal of Archaeological Science or Antiquity provide mainly traditional non-open access papers, with open access papers comprising only (respectively) 10%, less than 5% and less than 2% of all articles on the 1st of September 2024. These results are unlike observations made for articles dealing with machine learning applications in other disciplines such as soil science (Padarian et al., 2020), where paid-access routes vastly dominate that field of research. However, researchers have to keep in mind the existence of non-open access journal articles and book chapters to avoid a FUTON bias (Wentz, 2002), which will lead to not exploiting part of the wider research results in the field. This points to a wider adoption of both FAIR (Findable, Accessible, Interoperable, Reusable) data principles (Wilkinson et al., 2016; Nicholson et al., 2023) and CARE (Collective benefit, Authority to control, Responsibility, Ethics) principles (Carroll et al., 2020; Gupta et al., 2023). Nonetheless, if the degree of open access of machine learning applications in archaeology seems promising, the large disparity borne from colonialism remains unfortunately wide (Davis, 2020a). The difference in the number of publications from “Global North” and “Global South” countries can be directly linked with the gap in the economic development between higher and lower-income countries, leading to the underdevelopment of science and technology (Sagasti, 1973; Allik et al., 2020). Increasing collaborations between different countries and institutions could fill this gap between higher and lower-income countries (Sonnenwald, 2007).
@@ 4.2 How is machine learning used across archaeology?
In the earlier stage of machine learning applications in archaeology, many archaeological questions could only be answered by seeking experts in informatics, mathematics, and physics (Menze et al., 2006; Boon et al., 2009; Menze and Ur, 2012; Hörr et al., 2014). However a noticeable shift has occurred over time. Archaeologists have progressively embraced programming their own tools (e.g. Python, R, Matlab, Java), and have also turned to specialised software solutions such as Grilli and Remondino [2019], Garcia-Malsosa et al. [2021], Demjam et al. [2022b] or Casini et al. [2023]. Nowadays, in many studies, archaeologists are both asking and answering the questions themselves (Orengo and Garcia-Molsosa, 2019; Orengo et al., 2020; Yaworsky et al., 2020; Carter et al., 2021). The CAA Special Interest Group on Scientific Scripting Languages in Archaeology (SIG-SSLA ) is one example of this polyvalency of archaeologists developing scripting or machine learning tools for their problems. In certain instances, computer science specialists independently pose questions and generate tailored answers for specific applications (Toler-Franklin et al., 2010; Ushizima et al., 2020; Wunderlich et al., 2022). However, there is always a question of the ratio between the time spent developing new skills (Aldenderfer, 1998, pp. 109-112) and the increasing complexity, if not in programming, then in understanding, of machine learning methods, making it more and more difficult for an archaeologist to master all abilities needed to develop an entire workflow for a computer-based approach.
@@ 4.2.1 Classification of artefacts and animal remains
Classifying artefacts and animal (including human) remains into categories is a central aspect of archaeological work (Read, 2018). Archaeologists have looked for solutions in computer science since the second half of the twentieth century [cf. 1 Binford and Binford 1966, Orton 1980, pp. 156-178]. It is therefore not surprising to see machine learning applications developed for this purpose already very early [Fig. 5A]. Earlier papers generally applied methods such as decision trees, ensemble learning, Bayesian classifiers, as well as linear classifiers combined with nearest neighbour classifiers (Nguifo et al., 1997; Hörr et al., 2014; Mircea et al., 2015b; MacLeod, 2018; Canul-Ku et al., 2019; Pargeter et al., 2019; González-Molina et al., 2020). Since 2020, however, ANNs have gained more traction [e.g. Ushizima et al. 2020; Graham et al. 2020], and now represent the most commonly used family of methods applied for this task [Fig. 7, Ling et al. 2024]. ANNs, especially CNNs, are standard for computer vision tasks like object classification (Guo et al., 2016; Chai et al., 2021). Dimensionality reduction methods (e.g. PCA and MDA) are often performed alongside machine learning algorithms, and can be an effective way to improve model accuracy when applied correctly [e.g. Muzzall 2021]. However, their ability to distinguish categories of artefacts seems limited compared to machine learning methods such as random forest (Cole et al., 2022).
@@
Concerns about the validity of certain forms of typological classification, both human-made and computer-based, as a practice have been previously raised, especially from North American researchers (ada, 1991; Hörr et al., 2014). Other researchers, however, do not hesitate in using computational approaches for classification (Pargeter et al., 2019), and even creating new classification methodologies with the help of machine learning. Notably, Hörr and others, seeking a less subjective approach to typological classification, developed a three-layer method for this purpose [Hörr 2011, pp. 136-137, Hörr et al. 2014], successively incorporating unsupervised, semi-supervised, and supervised classification algorithms within a framework Fayyad et al. [1996] termed the Knowledge Discovery Process. This approach is promising for the analysis of archaeological material, and could prove useful for many archaeological studies. Such an approach invites critical questions on the artificiality of typological classifications and opens new avenues for future studies such as those dealing with ceramic material or stone tools. We can also mention Martín-Perea et al. [2020], who used a method similar to that of (Hörr et al., 2014) with a three-level pipeline to detect fossiliferous levels of an archaeological or palaeontological site.
@@
Assessing and comparing the outcomes of these diverse classifications is challenging, given the inherent variations in datasets, modelling techniques, and evaluation metrics employed across different studies. Nonetheless, various opinions have already emerged on the issue. Emmit et al. [2022] and Jalandoni [2022] are optimistic about using machine learning for the classification of archaeological material, though we could also add Hörr et al. [2014], who are more cautious but remain hopeful for the future of the practice. Others are less enthusiastic, however, and point to mixed results (Demján et al., 2022b; Lyons et al., 2022). Overall it seems likely that machine learning applications for the classification of archaeological material will spread not only within the scientific community but also among professionals of archaeology, conservation institutions, and a wider public. The ArchAIDE project is already a noteworthy example of an application of open-access machine learning tools for archaeology to serve the public at large (Gualandi et al., 2021; Anichini et al., 2021; Anichini and Gattiglia, 2022). Another example of a mobile application for ceramics classification is developed in Santos et al. [2024], and a broader project of automation and digitalisation tools for archaeological artefacts has been developed in the Automata initiative (Naso and Sciuto, 2025). The reconstruction of ceramics based only on few pottery sherds has also been achieved with machine learning methods (Stamatopoulos and Anagnostopoulos, 2016; Cardarelli, 2024). Another promising development has been made in Ruschioni et al. [2023], were they classified ceramics based not on their typology but rather on their chemical properties, which is more data-compatible with a machine learning process.
In summary, as one of the core elements of archaeological work, the use of machine learning for the classification of artefacts has been debated theoretically and tested since the early 2000s. While ceramics are the most represented type of artefact used for machine learning-based classification in our corpus, a wide variety of materials have also been used (e.g., ivory, stone tools, etc.). With the majority (62%, n = 18) of study cases using image data as inputs, the recent increase in performance of neural network models could spur further advances for artefact classification. The future of this discipline will also likely be linked with more data-driven material, standardised pre-processing, and user-friendly classification software, be it portable or not.